1. 600 J of work are done on a system in a process that decreases the system's thermal energy by 300 J .

How much heat energy is transferred?

2. An open container holds ice of mass 0.565 kg at a temperature of -16.6 ∘C . The mass of the container can be ignored. Heat is supplied to the container at the constant rate of 890 J/minute . The specific heat of ice to is 2100 J/kg⋅K and the heat of fusion for ice is 334×103J/kg.

How much time tmelts passes before the ice starts to melt?

From the time when the heating begins, how much time trise does it take before the temperature begins to rise above 0∘C?
Answer #1

Change of system's internal energy equals work done On it plus heat heat transferred to it.

∆U = W + Q

∆U = -300 J (since it has decreased)

W = 600 J

thus,

Q = ∆U - W = -300J - 600J = -900J

So 900J of heat are transferred. Negative sign indicates that heat is transferred from the system.